OpenStudy (anonymous):

MEDAL AND FAN PLEASE Find a(b) a= 4i + 3j, b= -4i + 4j -4 28 < -16, 12> < 0, 7>

OpenStudy (anonymous):

(4i+3j)(-4i+4j)

OpenStudy (anonymous):

??

OpenStudy (amoodarya):

do you mean ? dot product or cross product

OpenStudy (anonymous):

idk

OpenStudy (amoodarya):

\[(4i+3j).(-4i+4j)=4(-4) +3(4) =-16 +12=-4\] dot product is this . answer is a number

OpenStudy (anonymous):

thanks!@amoodarya

OpenStudy (amoodarya):

but if you want cross product ...answer is a vector

OpenStudy (amoodarya):

\[\begin{vmatrix} i & j &k \\ 4 &3 &0 \\ -4 &4 &0 \end{vmatrix}=\\0 i -0j +(16-(3(-4))k=(0,0,28)\]this is cross product